OpenStudy (anonymous):

Given that the diagonal of a square is (1-5*square root 2)/(4-4*square root2), calculate p^2, where p is the perimeter of this square. Given your answer in the form a+b*square root 2 where a and b are real numbers. I arrive at a ans of (51-10*square root 2)/(6-4*square root 2) but I am not sure about my answer.

OpenStudy (amistre64):

given that a square is just a 45-45-90 with sides of 1-1-sqrt(2) this seems to be a question related to that little phenomenon

OpenStudy (amistre64):

a side is p/4

OpenStudy (lgbasallote):

first of all...we know that P = 4s where s = side and the diagonal = s sqrt 2 so the relationship between perimeter and diagonal is... d = (p sqrt 2)/2 ...that's a good palce to start :D

OpenStudy (amistre64):

\[\frac{1-5\sqrt{2}}{4-4\sqrt{2}}=\sqrt{2}\] \[x=1\]

OpenStudy (amistre64):

so id say x; which is my side is:\[\dfrac{\frac{1-5\sqrt{2}}{4-4\sqrt{2}}}{\sqrt{2}}=x=side\]

OpenStudy (amistre64):

^4 that and "simplify" :)

OpenStudy (amistre64):

or rather; *4 that and then ^4 it ... :/

OpenStudy (amistre64):

ill get it eventually lol; *4 that and ^2 it

OpenStudy (amistre64):

\[\dfrac{\frac{1-5\sqrt{2}}{4-4\sqrt{2}}}{\sqrt{2}}\] \[\frac{1-5\sqrt{2}}{4\sqrt{2}-4(2)}\] \[\frac{1-5\sqrt{2}}{4(\sqrt{2}-2)}*4\] \[\frac{1-5\sqrt{2}}{\sqrt{2}-2}=p\]

OpenStudy (lgbasallote):

so... (1-5*square root 2)/(4-4*square root2) = (p sqrt2)/2 now we cancel 2... (1-5*square root 2)/(2-2*square root2) = (p sqrt2) divide both sides by sqrt 2 (1-5*square root 2)/(2sqrt 2-4) = p square both sides... p^2 = (1-5sqrt2)^2/(2sqrt2 - 4)^2 (1 -10sqrt2 + 100)/(8 -16sqrt2 + 16) = p^2

OpenStudy (amistre64):

\[\left(\frac{1-5\sqrt{2}}{\sqrt{2}-2}\right)^2=p^2\] \[\frac{1-10\sqrt{2}+250}{2-4\sqrt{2}+4}=p^2\] maybe

OpenStudy (lgbasallote):

lol...all three of us had different answers :)) this question is tricky

OpenStudy (amistre64):

i think he missed a "2" to begin with; "2"51-10sqrt(2) would make our answers match; not ours but mine and his :)

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=+1%2F8+%288%2B9+sqrt%282%29%29+-+%28%281-5sqrt%282%29%29%2F%284sqrt%282%29-4%282%29%29%29 wolfs simplification of : s=1/8 (8+9sqrt(2)) seems to zero out with mine; just saying ;)